# UltraVSR: Achieving Ultra-Realistic Video Super-Resolution with Efficient One-Step Diffusion Space

### v1

@@ Abstract.
Diffusion models have shown great potential in generating realistic image detail. However, adapting these models to video super-resolution (VSR) remains challenging due to their inherent stochasticity and lack of temporal modeling, which together hinder the preservation of temporal consistency across frames. Previous methods have attempted to mitigate this issue by incorporating motion information and temporal layers. However, motion estimation from low-resolution videos is often unreliable due to missing fine details. Furthermore, these approaches typically rely on deep temporal layers and require multiple sampling steps per frame, resulting in high computational costs and limiting their applicability to short video sequences. In this paper, we propose UltraVSR, a novel framework that enables ultra-realistic and temporal-coherent VSR through an efficient one-step diffusion space. A central component of UltraVSR is the Degradation-aware Restoration Schedule (DRS), which estimates a degradation factor from the low-resolution input and transforms iterative denoising process into a single-step reconstruction from from low-resolution to high-resolution videos. This design eliminates randomness from diffusion noise and significantly speeds up inference. To ensure temporal consistency, we propose a lightweight yet effective Recurrent Temporal Shift (RTS) module, composed of an RTS-convolution unit and an RTS-attention unit. By partially shifting feature components along the temporal dimension, these two units collaboratively facilitate effective feature propagation, fusion, and alignment across neighboring frames, without relying on explicit temporal layers. The RTS module is integrated into a pretrained text-to-image diffusion model and is further enhanced through Spatio-temporal Joint Distillation (SJD), which improves temporal coherence while preserving realistic details. Additionally, we introduce a Temporally Asynchronous Inference (TAI) strategy to capture long-range temporal dependencies under limited memory constraints. Extensive experiments show that UltraVSR achieves state-of-the-art performance, both qualitatively and quantitatively, in a single sampling step.
@@
ccs: Computing methodologies Reconstructionccs: Computing methodologies Image processingccs: Computing methodologies Computational photography
Figure 1. Quantitative and qualitative VSR comparisons. Top-left: An example on a real-world low-resolution (LR) video frame; Top-right: Quantitative comparison with state-of-the-art VSR approaches on the VideoLQ benchmark (Chan et al., 2022b), where our approach achieves leading scores across all three metrics with just a one-step (s1) sampling process; Bottom: Visual comparison across three consecutive frames. SISR approaches often suffer from temporal inconsistency, while VSR approaches typically lack fine texture details. In contrast, our approach produces visually realistic and temporally coherent results.
[…]
@@ 1. Introduction
Video super-resolution (VSR) aims to reconstruct high-resolution (HR) videos from low-resolution (LR) inputs, a task that has garnered significant attention in multimedia and computer vision communities. Unlike single-image super-resolution (SISR), VSR presents additional challenges due to the necessity of maintaining temporal consistency across frames while recovering high-frequency details. With the rapid advancement of deep learning, numerous VSR approaches (Chan et al., 2021, 2022a; Fuoli et al., 2019; Gupta et al., 2021; Jin et al., 2023; Leng et al., 2022; Xiao et al., 2024; Ai et al., 2024) have emerged over the past decade. However, despite notable progress, existing methods still struggle to restore realistic textures under complex and unknown real-world degradations.
@@
Figure 2. Comparison of runtime and metrics between diffusion-based VSR approaches. Note that the runtime is measured with the HR sequence upscaled to 720$\times$1280. Our approach achieves the fastest speed with better quality.
@@
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have achieved remarkable success in image synthesis (Rombach et al., 2022; Dhariwal and Nichol, 2021), image editing (Kawar et al., 2023; Yang et al., 2023) and SISR (Wang et al., 2023a, 2024b). However, adapting these diffusion models to VSR remains challenging due to the lack of temporal modeling capabilities. Additionally, the inherent randomness of the diffusion process further exacerbates temporal artifacts such as flickering across frames, e.g., the results of SUPIR (Yu et al., 2024) and OSEDiff (Wu et al., 2024) in Fig. 1. Recent approaches, such as MGLD-VSR (Yang et al., 2025) and Upscale-A-Video (Zhou et al., 2024), attempted to improve temporal consistency by integrating 3D convolutions, temporal attention mechanisms, and motion propagation modules into image diffusion models. Alternatively, text-to-video diffusion models have also been explored for VSR (Xie et al., 2025). Nevertheless, these approaches suffer from high computational costs due to the use of temporal layers and unreliable motion information estimated from LR videos. Moreover, these approaches require a large number of sampling steps for each frame, making inference inefficient and impractical, especially for memory-constrained processing of high-resolution and long-duration video sequences.
@@
This paper proposes UltraVSR, an efficient one-step diffusion model for real-world VSR. In contrast to multi-step VSR approaches, UltraVSR introduces a Degradation-aware Restoration Scheduling (DRS) that reformulates the conventional multi-step iterative denoising paradigm as a single-step reconstruction process. DRS first estimates a degradation factor from the input LR video and then maps it with the diffusion noise schedule, thereby achieving single-step reconstruction from LR input to HR video. This scheme eliminates the randomness in the diffusion and significantly improves computational efficiency. To maintain temporal consistency across frames, we develop a lightweight yet effective Recurrent Temporal Shift (RTS) module without explicit temporal layers. The RTS module enables efficient inter-frame feature propagation, fusion, and alignment through two novel units: (1) RTS-convolution unit, which partitions feature maps by channels and shifts groups temporally across adjacent frames for localized feature alignment using 2D convolution; (2) RTS-attention unit, which incorporates temporal shifting operations on $Key$/$Value$ tensors prior to self-attention computation, allowing the model to capture global contextual dependencies.
@@
  • We propose UltraVSR, an efficient one-step diffusion framework by introducing a Degradation-aware Restoration Scheduling (DRS) for high-quality video reconstruction. To the best of our knowledge, this is the first work to tackle the VSR problem in just one sampling step.
@@
  • We introduce Spatio-temporal Joint Distillation (SJD) to enhance temporal coherence and content realism, and propose Temporally Asynchronous Inference (TAI) for memory-efficient long-range dependency modeling.
  • Extensive experiments demonstrate that UltraVSR outperforms state-of-the-art methods in both visual quality and temporal consistency, while significantly accelerating inference speed.
@@ 2.1. Real-World VSR
Numerous VSR methods (Chan et al., 2021; Fuoli et al., 2019; Gupta et al., 2021; Li et al., 2023; Liang et al., 2022; Xiao et al., 2020; Leng et al., 2022) have been developed to enhance video quality under real-world degradation. For instance, Chan et al. proposed RealBasicVSR (Chan et al., 2022b), which removes degradations in videos by introducing a pre-cleaning module and a stochastic degradation scheme. More recently, Zhang et al. introduced a channel-attention-based recurrent model, RealViformer (Zhang and Yao, 2024), based on findings from investigating channel and spatial attention in a real-world setting. Despite these advancements, many approaches remain sensitive to complex, unknown degradations due to limited modeling capacity, leading to inadequate texture recovery and poor generalization.
@@ 2.2. Diffusion Model-based VSR
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated strong performance in generating realistic details for SISR (Wang et al., 2024c; Yue et al., 2024; Wang et al., 2024b; Wu et al., 2024; Wang et al., 2024a). For example, Yu et al. (Yu et al., 2024) propose SUPIR that expands the scope of image restoration by enhancing perceptual quality and enabling controlled generation via textual prompts, while Wu et al. (Wu et al., 2024) introduce OSEDiff, an efficient one-step super-resolution method based on variational score distillation (Yin et al., 2024; Wang et al., 2023b). Nevertheless, it is still a non-trivial challenge to scale up these models on VSR task due to the inherent randomness. Recent studies have attempted to address this issue in two ways. On one hand, Zhou et al.(Zhou et al., 2024) and Yang et al.(Yang et al., 2025) leverage image diffusion models and propose Upscale-A-Video and MGLD-VSR, respectively, by incorporating additional temporal layers and motion cues extracted from low-resolution (LR) videos. On the other hand, Xie et al.(Xie et al., 2025) build upon a pretrained text-to-video diffusion model and introduce STAR, which employs ControlNet(Zhang et al., 2023) to inject LR video information into the generation process. However, these approaches often rely on multiple sampling steps and complex temporal layers, resulting in high inference latency. To overcome these challenges, we propose an efficient one-step diffusion model for VSR.
@@ 3.1. Overview
During training, we incorporate LoRA layers (Hu et al., 2022) into both the VAE encoder and diffusion UNet, achieving efficient parameter tuning while preserving generation priors. To ensure temporal consistency throughout the reconstruction process from $V_{LR}$ to $V_{HR}$, we integrate the RTS module across multiple scales of both the diffusion UNet and VAE decoder, components identified as primary sources of temporal flickering artifacts in the original architecture (Yang et al., 2025; Zhou et al., 2024). SJD is further used to enforce that the one-step reconstruction model generates the HR video $V_{HR}$ with both spatial realism and temporal coherence. The detailed network structure of these models can be found in the supplementary materials. Furthermore, we also propose a Temporally Asynchronous Inference (TAI) strategy for memory-efficient long-range dependency modeling in Sec. 3.5.
@@ 3.3. Recurrent Temporal Shift Module
The key to VSR lies in effectively aggregating complementary spatio-temporal information from misaligned frames. To this end, we propose a lightweight yet effective Recurrent Temporal Shift (RTS) module that captures spatio-temporal dependencies without relying on computationally expensive temporal layers. As illustrated in Fig.3 (b), the RTS module performs temporal feature propagation by bidirectionally shifting partitioned features across adjacent frames in a parameter-free manner, inspired by(Lin et al., 2019; Cao et al., 2023; Munoz et al., 2021). It consists of two core components: the RTS-Convolution Unit and the RTS-Attention Unit, which collaborate to enable efficient information exchange and fusion across neighboring frames. To further improve computational efficiency, we apply channel-reducing and channel-increasing convolutions at the beginning and end of each unit, respectively.
@@ 3.4. Spatio-temporal Joint Distillation
where $\lambda$ balances content realism and temporal coherence of $V_{HR}$. This encourages the synthesized $V_{HR}$ toward higher perceptual realism and temporal coherence while suppressing visual artifacts and flickering. Specifically, the content-realistic gradient loss is formulated as:
@@
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, $z_{0}=\mathcal{E}(V_{HR})$, $V_{HR}=G_{\theta}(V_{LR})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $\mathcal{E}$ denotes the pretrained VAE encoder, $\omega(t)$ is a weighting coefficient to normalize the gradient magnitudes across varying noise levels (Yin et al., 2024). The temporal-consistency gradient loss enhances temporal coherence by penalizing discrepancies in the inter-frame gradients of two TRUs:
@@ 3.5. Temporally Asynchronous Inference
Figure 4. The previous inference scheme vs. the proposed Temporally Asynchronous Inference (TAI) scheme.
@@ 4.1. Experimental Settings
We implement UltraVSR using the PyTorch framework and the Diffusers (Patrick von Platen and Wolf, 2022) library. The input sequence length is set to 6, with the spatial size cropped to 512 $\times$ 512 resolution. We freeze the original spatial modules of the pretrained Stable Diffusion model to preserve its generation priors, while only training the newly introduced LoRA and RTS modules. Adam (Kingma and Ba, 2014) optimizer is employed as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We use a combination of MSE loss, perceptual loss, temporal GAN loss (Zhou et al., 2023), and SJD regularization loss to train UltraVSR. The real and fake TRUs are trained by minimizing a standard denoising objective (Ho et al., 2020; Vincent, 2011). The weighting coefficient $\lambda$ is set to 0.5. The initial learning rate is set to $5\times 10^{-5}$ for all models.
@@
We evaluate super-resolved video quality using three image-based quality metrics: MUSIQ (Ke et al., 2021), CLIP-IQA (Wang et al., 2023a), and MANIQA (Yang et al., 2022), along with one video-specific quality metric, DOVER (Wu et al., 2023). DOVER assesses both content quality and temporal consistency from both aesthetic and technical perspectives. In addition, the full-reference metrics (i.e., PSNR and SSIM) are also given as a reference. However, it should be noted that they only reflect certain aspects of performance and poorly respond to realistic visuals (Blau and Michaeli, 2018; Gu et al., 2022; Ledig et al., 2017; Yu et al., 2024; Jinjin et al., 2020), which can not reliably evaluate the performance of diffusion-based VSR models.
@@ 4.2. Evaluation with State-of-the-art Methods
To validate the efficacy of UltraVSR, we conduct both quantitative and qualitative comparisons with several state-of-the-art VSR methods: BasicVSR++ (Chan et al., 2022a), OVSR (Yi et al., 2021), RealBasicVSR (Chan et al., 2022b), VRT (Liang et al., 2024), RealViformer (Zhang and Yao, 2024), Upscale-A-Video (Zhou et al., 2024), MGLD-VSR (Shi et al., 2024), and STAR (Xie et al., 2025). For a more comprehensive evaluation, we also include comparisons with state-of-the-art SISR approaches (i.e., RealESRGAN (Wang et al., 2021), StableSR (Wang et al., 2024c), SUPIR (Yu et al., 2024), ResShift (Yue et al., 2024) and OSEDiff (Wu et al., 2024)) in the supplementary materials.
@@
As demonstrated in Tab. 1, we present the quantitative results with existing state-of-the-art VSR approaches. It can be seen that our UltraVSR consistently achieves the highest score across nearly all non-reference metrics on all synthetic benchmarks, demonstrating its superior performance. Although approaches like BasicVSR++ (Chan et al., 2022a) and VRT (Liang et al., 2024) have better performance in terms of PSNR or SSIM, they usually produce unsatisfactory results on real-world degraded videos (see Sec. 4.2.2), as reflected in their lower no-reference metric scores. As for real-world VSR dataset VideoLQ (Chan et al., 2022b), UltraVSR outperforms other methods across all non-reference metrics (MUSIQ: +1.94, CLIP-IQA: +0.0936, MANIQA: +0.0186, DOVER: +0.014). These notable improvements highlight UltraVSR’s strong ability to enhance complex and diverse real-world videos while preserving realistic details.
Table 2. Efficiency evaluations among diffusion-based VSR approaches. Note that the runtime is evaluated using a single RTX 8000 GPU. Both trainable/total parameters are measured for each model. The best results are highlighted.
@@
To further assess the effectiveness of the proposed UltraVSR, we present qualitative comparisons on both synthetic datasets (SPMCS (Yi et al., 2019) and YouHQ40 (Zhou et al., 2024) in Fig. 5) and real-world dataset (VideoLQ (Chan et al., 2022b) in Fig. 6). On synthetic videos, it can be observed that the competing approaches struggle to reconstruct accurate texture (e.g., the results of Upscale-A-Video (Zhou et al., 2024) and STAR (Xie et al., 2025) in the first row) and blurriness (e.g., the results of RealViformer (Zhang and Yao, 2024) and the result of Upscale-A-Video (Zhou et al., 2024) in the second row). In addition, MGLD-VSR (Shi et al., 2024) introduces color artifacts in the reconstructed frames. In contrast, our UltraVSR excels in restoring lost textures and structures under complex degradation scenarios, such as the outlines of statues and animal fur. On real-world videos, UltraVSR generates high-quality texture information that surpasses those of competing approaches. For instance, in sequence 004, UltraVSR not only recovers clear and realistic facial expressions but also reconstructs plausible exterior structure of the surrounding brick wall. In sequence 039, UltraVSR successfully restores intricate architectural details and decorations, demonstrating its strong robustness and generalization ability under challenging degradation scenarios. More comparisons can be found in the supplementary materials.
@@
Figure 7. Temporal consistency analysis across various approaches, including both SISR and VSR approaches, as well as different model variants.
@@
A key challenge in VSR is maintaining temporal consistency across frames. From a quantitative perspective, we compare the DOVER metric, an effective video quality assessment approach, against competing methods in Tab. 1. Due to the inherent randomness in the diffusion model, the DOVER scores of existing diffusion-based VSR methods are generally low. Benefiting from the proposed one-step diffusion space, our UltraVSR achieves the best score on all the datasets (e.g., +0.011 on the SPMCS dataset and +0.014 on the VideoLQ dataset). In addition, Fig. 7 presents visual comparisons against both SISR and VSR approaches. It can be found that due to the inherent randomness and insufficient temporal modeling, results from SUPIR (Yu et al., 2024) and OSEDiff (Wu et al., 2024) exhibit severe temporal inconsistencies. While existing VSR approaches alleviate this issue, they often sacrifice texture details. In contrast, our ultraVSR successfully reconstructs videos with both fine texture structure and temporal coherence.

### v2

@@ Abstract.
Diffusion models have shown great potential in generating realistic image detail. However, adapting these models to video super-resolution (VSR) remains challenging due to their inherent stochasticity and lack of temporal modeling. Previous methods have attempted to mitigate this issue by incorporating motion information and temporal layers. However, unreliable motion estimation from low-resolution videos and costly multiple sampling steps with deep temporal layers limit them to short sequences. In this paper, we propose UltraVSR, a novel framework that enables ultra-realistic and temporally-coherent VSR through an efficient one-step diffusion space. A central component of UltraVSR is the Degradation-aware Reconstruction Scheduling (DRS), which estimates a degradation factor from the low-resolution input and transforms the iterative denoising process into a single-step reconstruction from low-resolution to high-resolution videos. To ensure temporal consistency, we propose a lightweight Recurrent Temporal Shift (RTS) module, including an RTS-convolution unit and an RTS-attention unit. By partially shifting feature components along the temporal dimension, it enables effective propagation, fusion, and alignment across frames without explicit temporal layers. The RTS module is integrated into a pretrained text-to-image diffusion model and is further enhanced through Spatio-temporal Joint Distillation (SJD), which improves temporally coherence while preserving realistic details. Additionally, we introduce a Temporally Asynchronous Inference (TAI) strategy to capture long-range temporal dependencies under limited memory constraints. Extensive experiments show that UltraVSR achieves state-of-the-art performance, both qualitatively and quantitatively, in a single sampling step. Code is available at https://github.com/yongliuy/UltraVSR.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755117ccs: Computing methodologies Reconstructionccs: Computing methodologies Image processingccs: Computing methodologies Computational photography
Figure 1. Quantitative and qualitative VSR comparisons. Top-left: An example on a real-world low-resolution (LR) video frame; Top-right: Quantitative comparison with state-of-the-art VSR approaches on the VideoLQ benchmark (Chan et al., 2022b); Bottom: Visual comparison against both SISR approaches and VSR approaches across three consecutive frames. Our one-step (s1) approach achieves leading scores across all three metrics with visually realistic and temporally coherent results.
@@ 1. Introduction
Video super-resolution (VSR) aims to reconstruct high-resolution (HR) videos from low-resolution (LR) inputs, a task that has garnered significant attention in multimedia and computer vision communities. Unlike single-image super-resolution (SISR), VSR presents additional challenges due to the necessity of maintaining temporal consistency across frames while recovering high-frequency details. With the rapid advancement of deep learning, numerous VSR approaches (Chan et al., 2021, 2022a; Fuoli et al., 2019; Gupta et al., 2021; Jin et al., 2023; Leng et al., 2022; Xiao et al., 2024; Ai et al., 2024) have emerged over the past decade. However, these methods still struggle to restore realistic textures under complex and unknown real-world degradations.
@@
Figure 2. Comparison of runtime and metrics among diffusion-based VSR approaches. Note that the runtime is measured with the HR sequence upscaled to 720$\times$1280. Our approach achieves the fastest speed with better quality.
@@
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have achieved remarkable success in image synthesis (Rombach et al., 2022; Dhariwal and Nichol, 2021), image editing (Kawar et al., 2023; Yang et al., 2023) and SISR (Wang et al., 2023a, 2024b). However, adapting these models to VSR remains challenging due to the lack of temporal modeling capabilities. Additionally, the inherent randomness of the diffusion process further exacerbates temporal artifacts such as flickering across frames, e.g., the results of SUPIR (Yu et al., 2024) and OSEDiff (Wu et al., 2024) in Fig. 1. Recent approaches, such as MGLD-VSR (Yang et al., 2025) and Upscale-A-Video (Zhou et al., 2024), have attempted to improve temporal consistency by integrating 3D convolutions, temporal attention mechanisms, and motion propagation modules into image diffusion models. Alternatively, text-to-video diffusion models have also been explored for VSR (Xie et al., 2025). Nevertheless, these approaches suffer from high computational costs due to the use of temporal layers and unreliable motion information estimated from LR videos. Moreover, these approaches require a large number of sampling steps for each frame, making inference inefficient and impractical, especially for memory-constrained processing of high-resolution and long-duration video sequences.
@@
This paper proposes UltraVSR, an efficient one-step diffusion model for real-world VSR. In contrast to multi-step VSR approaches, UltraVSR introduces a Degradation-aware Reconstruction Scheduling (DRS) that reformulates the conventional multi-step iterative denoising paradigm as a single-step reconstruction process. DRS first estimates a degradation factor from the input LR video and then maps it with the diffusion noise schedule, thereby achieving single-step reconstruction from LR input to HR video. This scheme eliminates the randomness in the diffusion process and significantly improves computational efficiency. To maintain temporal consistency across frames, we develop a lightweight yet effective Recurrent Temporal Shift (RTS) module, which enables efficient inter-frame feature propagation, fusion, and alignment through two novel units: (1) RTS-convolution unit, which partitions feature maps by channels and shifts groups temporally across adjacent frames for localized feature alignment using 2D convolution; (2) RTS-attention unit, which incorporates temporal shifting operations on $Key$/$Value$ tensors prior to self-attention computation, allowing the model to capture global contextual dependencies.
@@
  • We propose UltraVSR, an efficient one-step diffusion framework by introducing a Degradation-aware Reconstruction Scheduling (DRS) for high-quality video reconstruction. To the best of our knowledge, this is the first work to tackle the VSR problem in just one sampling step.
@@
  • We introduce Spatio-temporal Joint Distillation (SJD) to enhance temporally coherence and content realism, and propose Temporally Asynchronous Inference (TAI) for memory-efficient long-range dependency modeling.
  • Extensive experiments show that UltraVSR outperforms state-of-the-art methods in both visual quality and temporal consistency, while significantly accelerating inference speed.
@@ 2.1. Real-World VSR
Numerous VSR methods (Chan et al., 2021; Fuoli et al., 2019; Gupta et al., 2021; Li et al., 2023; Liang et al., 2022; Xiao et al., 2020; Leng et al., 2022) have been developed to enhance video quality under real-world degradation. For instance, Chan et al. proposed RealBasicVSR (Chan et al., 2022b), which removes degradations in videos by introducing a pre-cleaning module and a stochastic degradation scheme. More recently, Zhang et al. introduced a channel-attention-based recurrent model, RealViformer (Zhang and Yao, 2024), based on findings from investigating channel and spatial attention in a real-world setting. Despite these advancements, many approaches remain sensitive to complex degradations due to limited modeling capacity, leading to poor generalization.
@@ 2.2. Diffusion Model-based VSR
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated strong performance in generating realistic details (Rombach et al., 2022; Kawar et al., 2023; Li et al., 2025; Yang et al., 2023, 2024; Blattmann et al., 2023). Recent studies have attempted to scale up diffusion models for the VSR task in two ways. On one hand, Zhou et al.(Zhou et al., 2024) and Yang et al.(Yang et al., 2025) leverage image diffusion models and propose Upscale-A-Video and MGLD-VSR, respectively, by incorporating additional temporal layers and motion information. On the other hand, Xie et al.(Xie et al., 2025) build upon a pretrained text-to-video diffusion model and introduce STAR, which employs ControlNet(Zhang et al., 2023) to inject LR video information into the generation process. However, these approaches often rely on multiple sampling steps and complex temporal layers, resulting in high inference latency. To overcome these challenges, we propose an efficient one-step diffusion model for real-world VSR.
@@ 3.1. Overview
During training, we incorporate LoRA layers (Hu et al., 2022) into both the VAE encoder and diffusion UNet, achieving efficient parameter tuning while preserving generation priors. To ensure temporal consistency throughout the reconstruction process from $V_{LR}$ to $V_{HR}$, we integrate the RTS module across multiple scales of both the diffusion UNet and VAE decoder, components identified as primary sources of temporal flickering artifacts in the original architecture (Yang et al., 2025; Zhou et al., 2024). SJD is further used to enforce that the one-step reconstruction model generates $V_{HR}$ with both spatial realism and temporally coherence. The detailed network architectures are provided in the supplementary materials. We also propose a Temporally Asynchronous Inference (TAI) strategy for memory-efficient long-range dependency modeling in Sec. 3.5.
@@ 3.3. Recurrent Temporal Shift Module
The key to VSR lies in effectively aggregating complementary spatio-temporal information from misaligned frames. To this end, we propose a lightweight yet effective Recurrent Temporal Shift (RTS) module that captures spatio-temporal dependencies without relying on computationally expensive temporal layers. As illustrated in Fig.3 (b), the RTS module performs temporal feature propagation by bidirectionally shifting partitioned features across adjacent frames in a parameter-free manner, inspired by(Lin et al., 2019; Cao et al., 2023; Munoz et al., 2021). It consists of two core components: the RTS-Convolution Unit and the RTS-Attention Unit, which collaborate to enable efficient information exchange across neighboring frames. To improve computational efficiency, we apply channel-reducing and channel-increasing convolutions at the beginning and end of each unit, respectively.
@@ 3.4. Spatio-temporal Joint Distillation
where $\lambda$ balances content realism and temporal coherence of $V_{HR}$. This encourages the synthesized $V_{HR}$ toward higher perceptual realism and temporally coherence while suppressing visual artifacts and flickering. The content-realistic gradient loss is formulated as:
@@
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, $z_{0}=\mathcal{E}(V_{HR})$, $V_{HR}=G_{\theta}(V_{LR})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $\mathcal{E}$ denotes the pretrained VAE encoder, $\omega(t)$ is a weighting coefficient to normalize the gradient magnitudes across varying noise levels (Yin et al., 2024). The temporal-consistency gradient loss enhances temporally coherence by penalizing discrepancies in the inter-frame gradients of two TRUs:
@@ 3.5. Temporally Asynchronous Inference
Figure 4. Previous inference scheme vs. the proposed TAI.
@@ 4.1. Experimental Settings
We implement UltraVSR using the PyTorch framework and the Diffusers (Patrick von Platen and Wolf, 2022) library. The input sequence length is set to 6, with the spatial size cropped to 512 $\times$ 512 resolution. We freeze the original spatial modules of the pretrained Stable Diffusion model while only training the newly introduced LoRA and RTS modules. Adam (Kingma and Ba, 2014) optimizer is employed as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We use a combination of MSE loss, perceptual loss, temporal GAN loss (Zhou et al., 2023), and SJD regularization loss to train UltraVSR. The real and fake TRUs are trained by minimizing a standard denoising objective (Ho et al., 2020; Vincent, 2011). The weighting coefficient $\lambda$ is set to 0.5. The learning rate is set to $5\times 10^{-5}$ for all models.
@@
We evaluate super-resolved video quality using three image-based quality metrics: MUSIQ (Ke et al., 2021), CLIP-IQA (Wang et al., 2023a), and MANIQA (Yang et al., 2022), along with one video-specific quality metric, DOVER (Wu et al., 2023). DOVER assesses both content quality and temporal consistency from both aesthetic and technical perspectives. In addition, the full-reference metrics (i.e., PSNR and SSIM) are also given as a reference. However, it should be noted that they poorly respond to realistic visuals (Blau and Michaeli, 2018; Gu et al., 2022; Ledig et al., 2017; Yu et al., 2024; Jinjin et al., 2020) and can not reliably evaluate the performance of diffusion-based VSR models.
@@ 4.2. Evaluation with State-of-the-art Methods
To validate the efficacy of UltraVSR, we conduct both quantitative and qualitative comparisons with state-of-the-art VSR methods: BasicVSR++ (Chan et al., 2022a), OVSR (Yi et al., 2021), RealBasicVSR (Chan et al., 2022b), VRT (Liang et al., 2024), RealViformer (Zhang and Yao, 2024), Upscale-A-Video (Zhou et al., 2024), MGLD-VSR (Shi et al., 2024), and STAR (Xie et al., 2025). Due to the extremely high memory requirements of STAR, we had to limit its output resolution to a maximum of 1440 pixels. In addition, we include comparisons with state-of-the-art SISR approaches (i.e., RealESRGAN (Wang et al., 2021), StableSR (Wang et al., 2024c), SUPIR (Yu et al., 2024), ResShift (Yue et al., 2024) and OSEDiff (Wu et al., 2024)) in the supplementary materials.
@@
As demonstrated in Tab. 1, we present the quantitative results with existing state-of-the-art VSR approaches. It can be seen that our UltraVSR consistently achieves the highest score across nearly all non-reference metrics on all synthetic benchmarks, demonstrating its superior performance. Although approaches like BasicVSR++ (Chan et al., 2022a) and VRT (Liang et al., 2024) have better performance in terms of PSNR or SSIM, they usually produce unsatisfactory results on real-world degraded videos (see Sec. 4.2.2), as reflected in their lower no-reference metric scores. As for the real-world VSR dataset VideoLQ (Chan et al., 2022b), UltraVSR outperforms other methods across all non-reference metrics (MUSIQ: +1.94, CLIP-IQA: +0.0936, MANIQA: +0.0186, DOVER: +0.014). These notable improvements highlight UltraVSR’s strong ability to enhance complex and diverse real-world videos while preserving realistic details.
Table 2. Efficiency evaluations among diffusion-based VSR approaches. Both trainable/total parameters are measured for each model. The best results are highlighted.
@@
To further assess the effectiveness of the proposed UltraVSR, we present qualitative comparisons on both synthetic datasets (SPMCS (Yi et al., 2019) and YouHQ40 (Zhou et al., 2024) in Fig. 5) and real-world dataset (VideoLQ (Chan et al., 2022b) in Fig. 6). On synthetic videos, it can be observed that the competing approaches struggle to reconstruct accurate texture (e.g., the results of Upscale-A-Video (Zhou et al., 2024) and STAR (Xie et al., 2025) in the first row) and blurriness (e.g., the results of RealViformer (Zhang and Yao, 2024) and the result of Upscale-A-Video (Zhou et al., 2024) in the second row). In addition, MGLD-VSR (Shi et al., 2024) introduces color artifacts in the reconstructed frames. In contrast, our UltraVSR excels in restoring lost textures and structures under complex degradation scenarios, such as the outlines of statues and animal fur. On real-world videos, UltraVSR demonstrates superior texture reconstruction capabilities compared to competing approaches. For instance, in sequence 004, it not only recovers clear and realistic facial expressions but also faithfully reconstructs the structural details of the surrounding brick wall. In sequence 039, UltraVSR successfully restores intricate architectural details and decorations, demonstrating its robustness and generalization ability under challenging degradation scenarios. More comparisons can be found in the supplementary materials.
@@
Figure 7. Temporal consistency analysis across SISR, VSR approaches, and different model variants.
@@
From a quantitative perspective, we compare the DOVER metric, an effective video quality assessment approach, against competing methods in Tab. 1. Due to the inherent randomness in the diffusion model, the DOVER scores of existing diffusion-based VSR methods are generally low. Benefiting from the proposed one-step diffusion space, our UltraVSR achieves the best score on all the datasets (e.g., +0.011 on the SPMCS dataset and +0.014 on the VideoLQ dataset). In addition, Fig. 7 presents visual comparisons against both SISR and VSR approaches. It can be found that due to the inherent randomness and insufficient temporal modeling, results from SUPIR (Yu et al., 2024) and OSEDiff (Wu et al., 2024) exhibit severe temporal inconsistencies. While existing VSR approaches alleviate this issue, they often sacrifice texture details. In contrast, our ultraVSR successfully reconstructs videos with both fine texture structure and temporal coherence.